OpenStudy (anonymous):

How do I solve this exponential equation 5(e^x-2)*e^3x-9=150?

OpenStudy (campbell_st):

well start by dividing both sides of the equation by 5 \[e^{x-2}\times e^{3x -9} = 30\] now apply the index law for multiplying the same base... add the powers which means \[e^{x -2 + 3x -9} = 30\] simplify that then take the base e log of both sides... this will allow you to solve for x. hope this helps

OpenStudy (anonymous):

I get .7522 as my answer.

OpenStudy (campbell_st):

nope... just doing a bit of mental arithmetic... the answer is closer to 5

OpenStudy (anonymous):

I got .9284 this time.

OpenStudy (campbell_st):

ok... taking the log of both sides \[\ln(e^{4x -11}) = \ln(30)\] or \[4x -11 = \ln(30)\] since \[\log_{a}(a^x) = x\] that law had to be applied here...

OpenStudy (anonymous):

Would you add 11 to 30 then divide by 4?

OpenStudy (campbell_st):

add 11 to ln(30) then divide the answer by 4

OpenStudy (anonymous):

3.600

OpenStudy (anonymous):

Is that closer to the answer?

OpenStudy (campbell_st):

thats a lot better

OpenStudy (anonymous):

Is that the correct answer?

OpenStudy (campbell_st):

well it is, given the way I worked it out...

OpenStudy (campbell_st):

you can check by substititing... is \[5 \times e^3.6 - 2) \times e^{3 \times 3.6 - 9} = 150...?\] you'll get an approximation because of rounding... see if it works

OpenStudy (campbell_st):

oops should read \[5 \times e^{3.6 -2} \times e^{3 \times 3.6 - 9} = 150...?\]